# Electric Dipole Moments in 5+3 Flavor Weak Effective Theory

### v1

@@ 5.4 Master formula for $d_{n}$
$C_{\widetilde{G}}$ $(5.4\pm 3.0)\cdot 10^{-22}$ $[C_{uG}]_{22}$ $(10.8\pm 6.1)\cdot 10^{-21}$ $[C_{dG}]_{33}$ $(7.7\pm 4.2)\cdot 10^{-22}$
$[C_{dG}]_{11}$ $(-3.7\pm 1.5)\cdot 10^{-17}$ $[C_{u\gamma}]_{22}$ $(-5.2\pm 2.5)\cdot 10^{-22}$ $[C_{d\gamma}]_{33}$ $(14.1\pm 7.0)\cdot 10^{-24}$
$[C_{dG}]_{22}$ $(15.1\pm 9.2)\cdot 10^{-21}$ $[C^{S1,RR}_{uu}]_{2222}$ $(-7.9\pm 4.0)\cdot 10^{-25}$ $[C^{V1,LR}_{uddu}]_{1331}$ $(6.0\pm 4.4)\cdot 10^{-23}$
$[C_{uG}]_{11}$ $(-18.3\pm 7.9)\cdot 10^{-18}$ $[C^{S8,RR}_{uu}]_{2222}$ $(17.9\pm 9.3)\cdot 10^{-26}$ $[C^{V8,LR}_{uddu}]_{1331}$ $(7.0\pm 2.5)\cdot 10^{-23}$
$[C_{d\gamma}]_{11}$ $(-89.0\pm 3.3)\cdot 10^{-18}$ $[C^{S1,RR}_{uu}]_{1122}$ $(-2.3\pm 1.2)\cdot 10^{-22}$ $[C^{V1,LR}_{dd}]_{1331}$ $(-19.7\pm 5.7)\cdot 10^{-23}$
$[C_{d\gamma}]_{22}$ $(3.0\pm 1.8)\cdot 10^{-19}$ $[C^{S8,RR}_{uu}]_{1122}$ $(-3.6\pm 4.0)\cdot 10^{-23}$ $[C^{V8,LR}_{dd}]_{1331}$ $(-52.0\pm 5.0)\cdot 10^{-23}$
$[C_{u\gamma}]_{11}$ $(23.5\pm 1.6)\cdot 10^{-18}$ $[C^{S1,RR}_{ud}]_{2211}$ $(-4.7\pm 2.4)\cdot 10^{-22}$ $[C^{V1,LR}_{dd}]_{2332}$ $(10.7\pm 5.8)\cdot 10^{-25}$
$[C^{V1,LR}_{uddu}]_{1111}$ $(-6.4\pm 3.4)\cdot 10^{-22}$ $[C^{S8,RR}_{ud}]_{2211}$ $(-28.2\pm 6.6)\cdot 10^{-23}$ $[C^{V8,LR}_{dd}]_{2332}$ $(14.3\pm 8.9)\cdot 10^{-25}$
$[C^{V8,LR}_{uddu}]_{1111}$ $(-10.3\pm 4.6)\cdot 10^{-22}$ $[C^{S1,RR}_{ud}]_{2222}$ $(3.0\pm 1.3)\cdot 10^{-26}$ $[C^{S1,RR}_{dd}]_{3333}$ $(-14.6\pm 7.7)\cdot 10^{-26}$
$[C^{V1,LR}_{uddu}]_{1221}$ $(-13.5\pm 6.4)\cdot 10^{-22}$ $[C^{S8,RR}_{ud}]_{2222}$ $(5.2\pm 3.1)\cdot 10^{-25}$ $[C^{S8,RR}_{dd}]_{3333}$ $(3.0\pm 1.7)\cdot 10^{-26}$
$[C^{V8,LR}_{uddu}]_{1221}$ $(-2.1\pm 1.1)\cdot 10^{-21}$ $[C^{S1,RR}_{uddu}]_{2112}$ $(25.9\pm 9.6)\cdot 10^{-22}$ $[C^{S1,RR}_{dd}]_{1133}$ $(-6.4\pm 3.3)\cdot 10^{-22}$
$[C^{V1,LR}_{dd}]_{1221}$ $(-7.0\pm 3.5)\cdot 10^{-22}$ $[C^{S8,RR}_{uddu}]_{2112}$ $(38.3\pm 9.3)\cdot 10^{-23}$ $[C^{S8,RR}_{dd}]_{1133}$ $(-0.9\pm 1.0)\cdot 10^{-22}$
$[C^{V8,LR}_{dd}]_{1221}$ $(-11.3\pm 5.6)\cdot 10^{-22}$ $[C^{S1,RR}_{uddu}]_{2222}$ $(-2.5\pm 1.6)\cdot 10^{-24}$ $[C^{S1,RR}_{dd}]_{2233}$ $(12.4\pm 7.8)\cdot 10^{-26}$
$[C^{S1,RR}_{uu}]_{1111}$ $(10.9\pm 4.8)\cdot 10^{-22}$ $[C^{S8,RR}_{uddu}]_{2222}$ $(-1.9\pm 1.3)\cdot 10^{-24}$ $[C^{S8,RR}_{dd}]_{2233}$ $(-3.7\pm 2.1)\cdot 10^{-25}$
$[C^{S8,RR}_{uu}]_{1111}$ $(-3.4\pm 1.8)\cdot 10^{-22}$ $[C^{S1,RR}_{uu}]_{1221}$ $(8.2\pm 5.0)\cdot 10^{-22}$ $[C^{S1,RR}_{dd}]_{1331}$ $(3.5\pm 2.0)\cdot 10^{-21}$
$[C^{S1,RR}_{dd}]_{1111}$ $(5.7\pm 2.9)\cdot 10^{-22}$ $[C^{S8,RR}_{uu}]_{1221}$ $(-24.9\pm 4.6)\cdot 10^{-23}$ $[C^{S8,RR}_{dd}]_{1331}$ $(-13.3\pm 2.5)\cdot 10^{-22}$
$[C^{S8,RR}_{dd}]_{1111}$ $(-17.8\pm 8.9)\cdot 10^{-23}$ $[C^{V1,LR}_{uddu}]_{2112}$ $(9.0\pm 1.4)\cdot 10^{-23}$ $[C^{S1,RR}_{dd}]_{2332}$ $(14.6\pm 8.0)\cdot 10^{-25}$
$[C^{S1,RR}_{dd}]_{2222}$ $(10.2\pm 6.3)\cdot 10^{-26}$ $[C^{V8,LR}_{uddu}]_{2112}$ $(21.2\pm 2.1)\cdot 10^{-23}$ $[C^{S8,RR}_{dd}]_{2332}$ $(2.9\pm 1.8)\cdot 10^{-24}$
$[C^{S8,RR}_{dd}]_{2222}$ $(12.1\pm 7.4)\cdot 10^{-26}$ $[C^{V1,LR}_{uddu}]_{2222}$ $(-2.1\pm 1.2)\cdot 10^{-25}$ $[C^{S1,RR}_{ud}]_{1133}$ $(-3.2\pm 1.4)\cdot 10^{-22}$
$[C^{S1,RR}_{dd}]_{1122}$ $(11.6\pm 8.1)\cdot 10^{-23}$ $[C^{V8,LR}_{uddu}]_{2222}$ $(-8.3\pm 5.3)\cdot 10^{-25}$ $[C^{S8,RR}_{ud}]_{1133}$ $(-13.6\pm 4.5)\cdot 10^{-23}$
$[C^{S8,RR}_{dd}]_{1122}$ $(-4.7\pm 2.1)\cdot 10^{-23}$ $[C^{V1,LR}_{uu}]_{1221}$ $(-26.4\pm 8.5)\cdot 10^{-24}$ $[C^{S1,RR}_{uddu}]_{1331}$ $(23.2\pm 8.1)\cdot 10^{-22}$
$[C^{S1,RR}_{dd}]_{1221}$ $(2.5\pm 1.0)\cdot 10^{-22}$ $[C^{V8,LR}_{uu}]_{1221}$ $(-73.9\pm 9.1)\cdot 10^{-24}$ $[C^{S8,RR}_{uddu}]_{1331}$ $(-0.6\pm 1.1)\cdot 10^{-22}$
$[C^{S8,RR}_{dd}]_{1221}$ $(-7.8\pm 2.6)\cdot 10^{-23}$ $[C^{S1,RR}_{ud}]_{2233}$ $(4.2\pm 2.3)\cdot 10^{-25}$
$[C^{S1,RR}_{ud}]_{1111}$ $(4.2\pm 2.3)\cdot 10^{-22}$ $[C^{S8,RR}_{ud}]_{2233}$ $(12.6\pm 6.8)\cdot 10^{-26}$
$[C^{S8,RR}_{ud}]_{1111}$ $(-12.6\pm 6.3)\cdot 10^{-23}$ $[C^{S1,RR}_{uddu}]_{2332}$ $(-20.0\pm 8.2)\cdot 10^{-25}$
$[C^{S1,RR}_{ud}]_{1122}$ $(2.6\pm 1.5)\cdot 10^{-22}$ $[C^{S8,RR}_{uddu}]_{2332}$ $(21.5\pm 9.6)\cdot 10^{-26}$
$[C^{S8,RR}_{ud}]_{1122}$ $(-8.7\pm 4.4)\cdot 10^{-23}$ $[C^{V1,LR}_{uddu}]_{2332}$ $(-5.3\pm 3.0)\cdot 10^{-26}$
$[C^{S1,RR}_{uddu}]_{1111}$ $(4.1\pm 2.2)\cdot 10^{-22}$ $[C^{V8,LR}_{uddu}]_{2332}$ $(3.5\pm 1.9)\cdot 10^{-26}$
$[C^{S8,RR}_{uddu}]_{1111}$ $(-12.5\pm 6.3)\cdot 10^{-23}$
$[C^{S1,RR}_{uddu}]_{1221}$ $(3.4\pm 1.5)\cdot 10^{-22}$
$[C^{S8,RR}_{uddu}]_{1221}$ $(-8.4\pm 4.3)\cdot 10^{-23}$
@@
Table 10: Master Formula for $d_{n}$ in WET in the JMS basis at the EW scale. The $\alpha_{I}^{n}$ have units of $e$ cm $\rm TeV^{2}$ and $e$ cm $\rm TeV$ for the 4-Fermi and dipole operators, respectively. The $90\%$ CL upper bound on $d_{n}$ is $1.8\times 10^{-26}$ $e$ cm.
@@ 6 Non-standard Higgs Couplings
This Lagrangian can arise in two theoretical frameworks that can be used for describing non-standard couplings of the Higgs to quarks, lepton, and gauge bosons. In SMEFT, gauge symmetry is linearly realized, and the Higgs boson belongs to a $SU_{L}(2)$ doublet and operators are organized according to their canonical dimension. In this framework, the SM Yukawa couplings arise at dim-4, they can always be chosen to be real and diagonal, and are determined by the quark and lepton masses. Non-standard Yukawa couplings arise at dim-6 and are given by the Lagrangian Buchmuller:1986zs ; Grzadkowski:2010es ; Hamoudou:2022tdn
@@
Also, the left and right-chiral fields are defined by $Q=(u_{L},d_{L})^{T},L=(\nu_{L},e_{L})^{T}$, $r=(u_{R},d_{R})^{T}$ and $\eta=(\nu_{R},e_{R})^{T}$ and $P_{\pm}=\frac{1}{2}\pm T_{3}$. $U$ is a matrix containing the Goldstone fields
$U=\textrm{exp}(2i\Phi/v),\qquad\Phi=\phi^{a}T^{a}=\frac{1}{\sqrt{2}}\begin{pmatrix}\phi^{0}&\phi^{+}\\ \phi^{-}&{-\phi^{0}}\end{pmatrix},$ (124)
@@ 6.1 Matching at the EW scale
Since EDMs can only constrain the imaginary parts of the WCs, they probe combinations of couplings such as
@@
At two-loop, Barr-Zee diagrams Barr:1990vd provide sizable contributions, and are sensitive to different combinations of Yukawa couplings. The diagrams are shown in Fig. 4. Integrating out the top-quark, Higgs boson, and $W$-boson at $\mu_{ew}$ leads to contributions to the quark and lepton dipole operators $C_{f\gamma}$ and to the quark chromo-dipole operator $C_{qG}$
$\textrm{Im}\left[C_{f\gamma}\right]_{ii}(\mu_{\rm ew})$ $=-12e\frac{\alpha}{(4\pi)^{3}}q_{f}q^{2}_{t}\frac{1}{m_{t}}\left[f(x_{t})\textrm{Re}(y^{f}_{ii})\textrm{Im}(y^{u}_{33})+g(x_{t})\textrm{Im}(y^{f}_{ii})\textrm{Re}(y^{u}_{33})\right]$
@@
$\textrm{Im}\left[C_{qG}\right]_{ii}(\mu_{\rm ew})$ $=2g_{s}\frac{\alpha_{s}}{(4\pi)^{3}}\frac{1}{m_{t}}\left[f(x_{t})\textrm{Re}(y^{q}_{ii})\textrm{Im}(y^{u}_{33})+g(x_{t})\textrm{Im}(y^{q}_{ii})\textrm{Re}(y^{u}_{33})\right],$ (131)

### v2

@@ 5.4 $d_{n}$
$C_{\widetilde{G}}$ $(5.5\pm 2.8)\cdot 10^{-22}$ $[C_{uG}]_{22}$ $(7.7\pm 3.9)\cdot 10^{-21}$ $[C_{dG}]_{33}$ $(1.1\pm 0.5)\cdot 10^{-21}$
$[C_{dG}]_{11}$ $(-3.7\pm 1.8)\cdot 10^{-17}$ $[C_{u\gamma}]_{22}$ $(-7.4\pm 3.4)\cdot 10^{-22}$ $[C_{d\gamma}]_{33}$ $(2.0\pm 1.1)\cdot 10^{-23}$
$[C_{dG}]_{22}$ $(1.5\pm 0.9)\cdot 10^{-20}$ $[C^{S1,RR}_{uu}]_{2222}$ $(-5.6\pm 3.0)\cdot 10^{-25}$ $[C^{V1,LR}_{uddu}]_{1331}$ $(6.0\pm 3.1)\cdot 10^{-22}$
$[C_{uG}]_{11}$ $(-1.9\pm 0.9)\cdot 10^{-17}$ $[C^{S8,RR}_{uu}]_{2222}$ $(1.3\pm 0.5)\cdot 10^{-25}$ $[C^{V8,LR}_{uddu}]_{1331}$ $(1.5\pm 0.9)\cdot 10^{-22}$
$[C_{d\gamma}]_{11}$ $(-8.9\pm 0.4)\cdot 10^{-17}$ $[C^{S1,RR}_{uu}]_{1122}$ $(-2.3\pm 1.1)\cdot 10^{-22}$ $[C^{V1,LR}_{dd}]_{1331}$ $(8.1\pm 6.4)\cdot 10^{-22}$
$[C_{d\gamma}]_{22}$ $(3.0\pm 1.7)\cdot 10^{-19}$ $[C^{S8,RR}_{uu}]_{1122}$ $(-3.6\pm 4.1)\cdot 10^{-23}$ $[C^{V8,LR}_{dd}]_{1331}$ $(-1.5\pm 0.2)\cdot 10^{-21}$
$[C_{u\gamma}]_{11}$ $(2.5\pm 0.2)\cdot 10^{-17}$ $[C^{S1,RR}_{ud}]_{2211}$ $(-4.8\pm 2.6)\cdot 10^{-22}$ $[C^{V1,LR}_{dd}]_{2332}$ $(1.1\pm 0.6)\cdot 10^{-24}$
$[C^{V1,LR}_{uddu}]_{1111}$ $(-4.3\pm 2.4)\cdot 10^{-22}$ $[C^{S8,RR}_{ud}]_{2211}$ $(-2.8\pm 0.7)\cdot 10^{-22}$ $[C^{V8,LR}_{dd}]_{2332}$ $(4.1\pm 2.5)\cdot 10^{-24}$
$[C^{V8,LR}_{uddu}]_{1111}$ $(-6.6\pm 3.7)\cdot 10^{-22}$ $[C^{S1,RR}_{ud}]_{2222}$ $(2.9\pm 1.4)\cdot 10^{-26}$ $[C^{S1,RR}_{dd}]_{3333}$ $(-2.2\pm 1.1)\cdot 10^{-25}$
$[C^{V1,LR}_{uddu}]_{1221}$ $(-9.0\pm 4.7)\cdot 10^{-22}$ $[C^{S8,RR}_{ud}]_{2222}$ $(5.2\pm 3.3)\cdot 10^{-25}$ $[C^{S8,RR}_{dd}]_{3333}$ $(4.6\pm 2.5)\cdot 10^{-26}$
$[C^{V8,LR}_{uddu}]_{1221}$ $(-1.4\pm 0.7)\cdot 10^{-21}$ $[C^{S1,RR}_{uddu}]_{2112}$ $(2.6\pm 1.0)\cdot 10^{-21}$ $[C^{S1,RR}_{dd}]_{1133}$ $(-6.4\pm 3.2)\cdot 10^{-22}$
$[C^{V1,LR}_{dd}]_{1221}$ $(-4.8\pm 2.6)\cdot 10^{-22}$ $[C^{S8,RR}_{uddu}]_{2112}$ $(3.8\pm 1.0)\cdot 10^{-22}$ $[C^{S8,RR}_{dd}]_{1133}$ $(-9.7\pm 9.5)\cdot 10^{-23}$
$[C^{V8,LR}_{dd}]_{1221}$ $(-7.4\pm 3.1)\cdot 10^{-22}$ $[C^{S1,RR}_{uddu}]_{2222}$ $(-2.5\pm 1.5)\cdot 10^{-24}$ $[C^{S1,RR}_{dd}]_{2233}$ $(1.3\pm 0.7)\cdot 10^{-25}$
$[C^{S1,RR}_{uu}]_{1111}$ $(1.1\pm 0.5)\cdot 10^{-21}$ $[C^{S8,RR}_{uddu}]_{2222}$ $(-1.9\pm 1.3)\cdot 10^{-24}$ $[C^{S8,RR}_{dd}]_{2233}$ $(-3.7\pm 2.1)\cdot 10^{-25}$
$[C^{S8,RR}_{uu}]_{1111}$ $(-3.4\pm 1.5)\cdot 10^{-22}$ $[C^{S1,RR}_{uu}]_{1221}$ $(8.3\pm 4.2)\cdot 10^{-22}$ $[C^{S1,RR}_{dd}]_{1331}$ $(3.5\pm 2.0)\cdot 10^{-21}$
$[C^{S1,RR}_{dd}]_{1111}$ $(5.7\pm 3.1)\cdot 10^{-22}$ $[C^{S8,RR}_{uu}]_{1221}$ $(-2.5\pm 0.4)\cdot 10^{-22}$ $[C^{S8,RR}_{dd}]_{1331}$ $(-1.3\pm 0.3)\cdot 10^{-21}$
$[C^{S8,RR}_{dd}]_{1111}$ $(-1.8\pm 0.9)\cdot 10^{-22}$ $[C^{V1,LR}_{uddu}]_{2112}$ $(5.4\pm 1.4)\cdot 10^{-22}$ $[C^{S1,RR}_{dd}]_{2332}$ $(1.4\pm 0.8)\cdot 10^{-24}$
$[C^{S1,RR}_{dd}]_{2222}$ $(1.0\pm 0.6)\cdot 10^{-25}$ $[C^{V8,LR}_{uddu}]_{2112}$ $(7.2\pm 0.5)\cdot 10^{-22}$ $[C^{S8,RR}_{dd}]_{2332}$ $(2.9\pm 1.6)\cdot 10^{-24}$
$[C^{S8,RR}_{dd}]_{2222}$ $(1.2\pm 0.6)\cdot 10^{-25}$ $[C^{V1,LR}_{uddu}]_{2222}$ $(-8.1\pm 4.5)\cdot 10^{-25}$ $[C^{S1,RR}_{ud}]_{1133}$ $(-3.2\pm 1.4)\cdot 10^{-22}$
$[C^{S1,RR}_{dd}]_{1122}$ $(1.2\pm 0.7)\cdot 10^{-22}$ $[C^{V8,LR}_{uddu}]_{2222}$ $(-2.8\pm 1.8)\cdot 10^{-24}$ $[C^{S8,RR}_{ud}]_{1133}$ $(-1.4\pm 0.5)\cdot 10^{-22}$
$[C^{S8,RR}_{dd}]_{1122}$ $(-4.7\pm 2.4)\cdot 10^{-23}$ $[C^{V1,LR}_{uu}]_{1221}$ $(6.3\pm 7.9)\cdot 10^{-23}$ $[C^{S1,RR}_{uddu}]_{1331}$ $(2.3\pm 1.0)\cdot 10^{-21}$
$[C^{S1,RR}_{dd}]_{1221}$ $(2.5\pm 1.0)\cdot 10^{-22}$ $[C^{V8,LR}_{uu}]_{1221}$ $(-2.6\pm 0.3)\cdot 10^{-22}$ $[C^{S8,RR}_{uddu}]_{1331}$ $(-5.8\pm 11.6)\cdot 10^{-23}$
$[C^{S8,RR}_{dd}]_{1221}$ $(-7.8\pm 2.0)\cdot 10^{-23}$ $[C^{S1,RR}_{ud}]_{2233}$ $(1.5\pm 0.7)\cdot 10^{-25}$
$[C^{S1,RR}_{ud}]_{1111}$ $(4.2\pm 2.1)\cdot 10^{-22}$ $[C^{S8,RR}_{ud}]_{2233}$ $(4.9\pm 2.0)\cdot 10^{-26}$
$[C^{S8,RR}_{ud}]_{1111}$ $(-1.3\pm 0.7)\cdot 10^{-22}$ $[C^{S1,RR}_{uddu}]_{2332}$ $(-9.7\pm 4.9)\cdot 10^{-25}$
$[C^{S1,RR}_{ud}]_{1122}$ $(2.6\pm 1.4)\cdot 10^{-22}$ $[C^{S8,RR}_{uddu}]_{2332}$ $(1.2\pm 0.6)\cdot 10^{-25}$
$[C^{S8,RR}_{ud}]_{1122}$ $(-8.8\pm 4.4)\cdot 10^{-23}$ $[C^{V1,LR}_{uddu}]_{2332}$ $(-2.7\pm 1.5)\cdot 10^{-25}$
$[C^{S1,RR}_{uddu}]_{1111}$ $(4.2\pm 2.2)\cdot 10^{-22}$ $[C^{V8,LR}_{uddu}]_{2332}$ $(7.5\pm 3.6)\cdot 10^{-26}$
$[C^{S8,RR}_{uddu}]_{1111}$ $(-1.3\pm 0.7)\cdot 10^{-22}$
$[C^{S1,RR}_{uddu}]_{1221}$ $(3.4\pm 1.4)\cdot 10^{-22}$
$[C^{S8,RR}_{uddu}]_{1221}$ $(-8.4\pm 3.9)\cdot 10^{-23}$
@@
Table 10: Master formula for $d_{n}$ in WET in the JMS basis at the EW scale. The $\alpha_{I}^{n}$ have units of $e$ cm $\rm TeV^{2}$ and $e$ cm $\rm TeV$ for the four-fermion and dipole operators, respectively. The $90\%$ CL upper bound on $d_{n}$ is $1.8\times 10^{-26}$ $e$ cm.
@@ 6 Non-standard Higgs Couplings
This Lagrangian can arise in two theoretical frameworks that can be used for describing non-standard couplings of the Higgs to quarks, lepton, and gauge bosons. In SMEFT, gauge symmetry is linearly realized, the Higgs boson belongs to a $SU_{L}(2)$ doublet and operators are organized according to their canonical dimension. In this framework, the SM Yukawa couplings arise at dim-4, they can always be chosen to be real and diagonal, and are determined by the quark and lepton masses. Non-standard Yukawa couplings arise at dim-6 and are given by the Lagrangian Buchmuller:1986zs ; Grzadkowski:2010es ; Hamoudou:2022tdn
@@
Also, the left and right-chiral fields are defined by $Q=(u_{L},d_{L})^{T},L=(\nu_{L},e_{L})^{T}$, $r=(u_{R},d_{R})^{T}$ and $\eta=(\nu_{R},e_{R})^{T}$ and $P_{\pm}=(1\pm\tau_{3})/2$, with $\bm{\tau}$ Pauli matrices. $U$ is a matrix containing the Goldstone fields
$U=\textrm{exp}(2i\Phi/v),\qquad\Phi=\phi^{a}\frac{\tau^{a}}{2}=\frac{1}{\sqrt{2}}\begin{pmatrix}\frac{\phi^{0}}{\sqrt{2}}&\phi^{+}\\ \phi^{-}&{-\frac{\phi^{0}}{\sqrt{2}}}\end{pmatrix},$ (124)
@@ 6.1 Matching at the EW scale
Here the matching scale $\mu_{ew}$ can be set to $m_{h}$. Since EDMs can only constrain the imaginary parts of the WCs, they probe combinations of couplings such as
@@
At two loops, Barr-Zee diagrams Barr:1990vd provide sizable contributions, and are sensitive to different combinations of Yukawa couplings. The diagrams are shown in Fig. 4. Integrating out the top-quark, Higgs boson, and $W$-boson at $\mu_{ew}$ leads to contributions to the quark and lepton dipole operators $C_{f\gamma}$ and to the quark chromo-dipole operator $C_{qG}$
$\textrm{Im}\left[C_{f\gamma}\right]_{ii}(\mu_{ew})$ $=-12e\frac{\alpha}{(4\pi)^{3}}q_{f}q^{2}_{t}\frac{1}{m_{t}}\left[f(x_{t})\textrm{Re}(y^{f}_{ii})\textrm{Im}(y^{u}_{33})+g(x_{t})\textrm{Im}(y^{f}_{ii})\textrm{Re}(y^{u}_{33})\right]$
@@
$\textrm{Im}\left[C_{qG}\right]_{ii}(\mu_{ew})$ $=2g_{s}\frac{\alpha_{s}}{(4\pi)^{3}}\frac{1}{m_{t}}\left[f(x_{t})\textrm{Re}(y^{q}_{ii})\textrm{Im}(y^{u}_{33})+g(x_{t})\textrm{Im}(y^{q}_{ii})\textrm{Re}(y^{u}_{33})\right],$ (131)